# Bayesian mixture modeling using a mixture of finite mixtures with normalized inverse Gaussian weights

### v1

@@ 1 Introduction
$f(y|\theta_{1},\dots,\theta_{M},\pi)=\sum_{j=1}^{M}\pi_{j}f(y|\theta_{j}),$ (1.1)
@@
where $M\in\{1,2,\dots\}$ is the number of components, $y\in\mathbb{R}^{d}$ is a $d$-dimensional observation vector, and $\pi=(\pi_{1},\dots,\pi_{M})^{\top}$ is the mixing weight vector, satisfying $\sum_{j=1}^{M}\pi_{j}=1$ and $\pi_{j}>0$. For $j=1,\dots,M$, $f(\cdot|\theta_{j})$ is a probability density or probability mass function parameterized by the component-specific parameter $\theta_{j}$. The function $f(\cdot|\theta_{j})$ is also called the kernel of the mixture model. Each observation is assumed to arise from one of the $M$ components, and each component is weighted based on its frequency of occurrence. Mixture models are important statistical models used in model-based clustering and density estimation because they can model complex data-generating distributions in random phenomena by using multiple components. Finite mixture models have been applied in many fields, for instance, sociology (Handcock et al., 2007), economics (Frühwirth-Schnatter and Kaufmann, 2008) and genetics (McLachlan et al., 2002) and so on. In application, the determination of $M$ is a very important issue. Correctly estimating $M$ improves model interpretability, the accuracy of kernel parameter estimates and model predictions, and reduces computation time. Although various methods have been proposed for the selection of $M$, such as model selection criteria and hypothesis testing, using these criterion makes it difficult to quantify the uncertainty of $M$, and also introduces the bias that model selection is carried out. The comprehensive review of (finite) mixture models is provided by McLachlan (2000); Frühwirth-Schnatter (2006); Frühwirth-Schnatter et al. (2019); McLachlan et al. (2019). In this paper, we focus on the case where $M$ is finite and make a clear distinction between the number of components $M$ and the number of clusters $k$. In general, the number of components is written as $M=k+M_{na}$, where $k$ is the number of components for which data are actually assigned, and $M_{na}$ is the number of empty components (see also Argiento and De Iorio, 2022).
@@
In Bayesian analysis, a finite mixture model that assumes prior distributions for the mixing weights and the number of components is usually employed (see, e.g. Nobile, 1994; Miller and Harrison, 2018). Such a model is also called a mixture of finite mixtures (MFM), and the model is often used, as well as infinite mixture models represented by Dirichlet process mixture models. Although the reversible jump algorithm (Green, 1995; Richardson and Green, 1997) has been used to obtain samples from the posterior distribution based on MFM, it faces significant computational and implementation challenges. Methods based on marginal likelihoods have also been proposed (see, e.g. Nobile and Fearnside, 2007), while, as with reversible jump, the computational aspect is an issue. A method called sparse MFM (Rousseau and Mengersen, 2011; Malsiner-Walli et al., 2016) has also been proposed, which focuses on estimating the number of clusters $k$ by using the overfitting model with a large fixed $M$, and choosing the prior distribution of the component rate well. Under same conditions, they showed that $k$ has consistency for true $M$. Since it allows for many empty components, we cannot estimate the number of components $M$ and it is not easy to quantify the uncertainty. Recently, the use of nonparametric Bayesian methods for MFM to estimate $M$ has attracted much attention. In Miller and Harrison (2018), the exchangeable partition distribution is derived by marginalizing out $M$, and the restaurant process is constructed to overcome computational difficulties. In Frühwirth-Schnatter et al. (2021), a generalized MFM in which the parameters of the Dirichlet distribution depend on $M$ is proposed. Miller and Harrison (2018) derived the theoretical result that the marginal posterior probabilities of $k$ and $M$ are asymptotically equivalent when the sample size is large, while only a sampling of $k$ is obtained as with general nonparametric Bayesian methods. Thus, the number of components $M$ cannot be sampled directly from the posterior distribution. Although Frühwirth-Schnatter et al. (2021)’s method can sample $M$ directly, as a sparse MFM, it has the disadvantage of producing a large number of empty components and overestimating $M$. In addition, most MFM studies use a symmetric Dirichlet distribution as the mixing weight vector, and the choice of a hyperparameter is very important. Although it is possible to set a prior distribution and learn from the data, the Metropolis-Hastings algorithm is required.
@@
The main purposes of this paper are 1) sampling $M$ directly and efficiently compared to MFM based on the Dirichlet distribution; 2) estimating $M$ reasonably by suppressing $M_{na}$; 3) proposing a MFM that is robust to the choice of hyperparameters of the distribution on a simplex. To this end, we employ a normalized inverse Gaussian distribution as the mixing weight vector of the finite mixture model. To the best of our knowledge, there are few studies using the normalized inverse Gaussian distributions as the mixing weights in the field of finite mixture models. In Bayesian nonparametric inference with infinite mixtures, Lijoi et al. (2005) proposed a normalized inverse Gaussian process and showed some analytical results of the posterior distribution under the process. Similarly, since Miller and Harrison (2018) is a finite version of the Dirichlet process, the model presented in this paper corresponds to a kind of finite alternative of Lijoi et al. (2005). Furthermore, leveraging data augmentation with a latent gamma random variable and the result of Argiento and De Iorio (2022), we construct an efficient posterior sampling algorithm.
@@
This paper is organized as follows. In Section 2, we introduce the MFM and its equivalent representation using discrete probability measures. We also discuss the data augmentation and the conditional posterior distribution of $M_{na}$, which are crucial for constructing the computational algorithm.In Section 3, we present the proposal method as well as the posterior computation algorithm. In Section 4, we illustrate some numerical studies to compare the proposed method with existing methods. R code implementing the proposed methods is available at Github repository:
@@ 2 Mixture of finite mixtures
We introduce a mixture of finite mixtures and its equivalent representation of the discrete probability measure. The proposed model and the posterior computation algorithm presented in the next section are largely based on the basic model presented in this section.
@@ 2.1 Formulation
Let each observations $Y_{1},\dots,Y_{n}$ be univariate or multivariate in an Euclidean space. A MFM model is a statistical model defined by the following hierarchical representation.
@@
$\begin{split}&M-1\sim q_{M},\quad q_{M}\ \text{is a probability mass function (p.m.f.) on}\ \{0,1,2,\dots\},\\ &\pi|M\sim P_{\pi}(\pi|M),\\ &c_{i}|M,\pi\sim\mathrm{Categorical}_{M}(\pi),\\ &\xi_{m}|M\overset{\mathrm{i.i.d}}{\sim}p_{0}(\xi),\quad m=1,\dots,M,\\ &\theta_{i}|c_{i},\xi\overset{\mathrm{ind}}{\sim}\delta_{\xi_{c_{i}}}(\theta_{i}),\quad i=1,\dots n,\\ &Y_{i}|\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}|\theta_{i}),\quad i=1,\dots,n,\end{split}$ (2.1)
@@
where $f(y|\theta_{i})$ is a parametric model with parameter $\theta_{i}$, $\delta_{x}(\cdot)$ is a point mass at $x$ and $\xi$ is a element of a vector $\xi=(\xi_{1},\dots,\xi_{M})^{\top}$. $p_{0}(\xi)$ is a prior density function of the parameter in mixture components. The parameter space is denoted by $\Theta\subset\mathbb{R}^{d}$. Each $c_{i}$ is a latent allocation indicates to which component each observation is allocated. Given the number of components $M$, $P_{\pi}(\pi|M)$ is a probability distribution on the $M$-dimensional unit simplex, in that, this is the prior distribution for the mixing weights. We focus the case where the mixing weights can be further hierarchically expressed as follows:
@@
$\begin{split}&S_{m}|M\overset{\mathrm{ind}}{\sim}h,\quad h\ \text{is a probability density function (p.d.f.) on}\ (0,\infty),\\ &\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim P_{\pi}(\pi|M),\quad T=\sum_{m=1}^{M}S_{m},\end{split}$ (2.2)
where $S_{1},\dots,S_{M}$ is unnormalized weights and $\pi$ is (normalized) mixing weight vector. This representation is the most basic way to generate a probability distribution on the $d$-dimensional unit simplex $\mathbb{S}_{d}:=\{w=(w_{1},\dots,w_{d});w_{i}\geq 0,\sum_{i=1}^{d}w_{i}=1\}$. One of the most famous examples of a distribution on $\mathbb{S}_{d}$ is the Dirichlet distribution. If $S_{m}|M\sim\text{Gamma}(\gamma_{i},1)$ in (2.2), $\pi\sim\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$, where $\text{Gamma}(a,b)$ is the gamma distribution with shape parameter $a>0$ and scale parameter $\beta>0$, and $\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$ is the Dirichlet distribution with parameter $\gamma=(\gamma_{1},\dots,\gamma_{d})>0$. If $\gamma_{1}=\cdots=\gamma_{d}=:\gamma$, the distribution obtained by normalization is $\text{Dirichlet}(\gamma,\dots,\gamma)$. The symmetric Dirichlet distribution is often used in the framework of MFM, because of conjugacy with categorical distributions and simplicity of computation. The symmetric structure of $\text{Dirichlet}(\gamma,\dots,\gamma)$ is also essential for marginalizing out $M$ and deriving the exchangeable partition distribution. However, it is known that the estimation result is sensitive to the choice of the shape parameter $\gamma$. For example, Miller and Harrison (2018) recommended to use $\gamma=1$ as a default choice and the value works well in many cases. The use of a small $\gamma$ was recommended in terms of sparse or generalized MFMs (Rousseau and Mengersen, 2011; Malsiner-Walli et al., 2016; Frühwirth-Schnatter et al., 2021). A small value of $\gamma$ indicates that many empty components can be created. Although most previous studies allow for the appearance of the empty components, there are several problems. First, $M$ tends to be overestimated, making the interpretation of the model difficult. Second, the existing of the empty components decreases the predictive performance of the model, as we will see in a later section through density estimations. Finally, they lead to an increase in computation time. When we focus on the estimation of the number of components $M$, it is desirable to have few empty components. In other words, the discrepancy between $M$ and $k$ should be small. To achieve this, we use the normalized inverse Gaussian distribution as mixing weights, instead of the Dirichlet distribution.
@@ 2.2 Equivalent representations using discrete probability measures
We give an equivalent representation of the MFM. We can construct a discrete measure $P(\cdot)=\sum_{m=1}^{M}\pi_{m}\delta_{\xi_{m}}(\cdot)$ in the parameter space $\Theta$, almost surely, where $M$, $\xi$ and $\pi$ are realizations from the distributions $q_{M}$, $p_{0}$ and $P_{\pi}$, respectively. Let $\theta_{1},\dots,\theta_{n}$ be random variables according to $P$. Then, the model (2.1) is equivalent to the following hierarchical representation using a random measure $P$.
$\begin{split}&Y_{i}|\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}|\theta_{i}),\quad i=1,\dots,n,\\ &\theta_{1},\dots,\theta_{n}|P\overset{\mathrm{i.i.d}}{\sim}P,\\ &P\sim\mathcal{P}(q_{M},h,p_{0}),\end{split}$ (2.3)
where $\mathcal{P}$ is a probability distribution of $P$ with parameters $q_{M}$, $h$ and $p_{0}$. The representation (2.3) is the MFM described as a nonparametric Bayesian framework with infinite mixtures. If we replace $\mathcal{P}$ with the Dirichlet process, (2.3) represents the well-known Dirichlet process mixture models (e.g., Escobar and West, 1995). If we replace $\mathcal{P}$ with the normalized inverse Gaussian process, (2.3) represents the normalized inverse Gaussian process mixture models (Lijoi et al., 2005). For MFM, Argiento and De Iorio (2022) proposed a normalized independent finite point process (Norm-IFPP), which is a class of flexible prior distributions for $P$. We employ the representation (2.3) with the Norm-IFPP. The advantages are as follows. We can directly estimate $M$ and $k$. This is a major difference from Miller and Harrison (2018), which estimates $M$ indirectly through the number of clusters $k$. Furthermore, an efficient Gibbs sampler can be constructed by incorporating the data augmentation with a latent Gamma random variable. This data augmentation enables us to overcome the lack of conjugacy in the categorical distribution. Instead of using the probability density function of $P_{\pi}$, we can use the density function of $h$. This is the key to building an efficient MCMC algorithm for the proposed model.
Our study is in the spirit of nonparametric Bayes in that it accounts for uncertainty in the prior distribution of the parameters by means of a random measure based on the Norm-IFPP. The details of Norm-IFPP and and the independent finite point process (IFPP) are given in Argiento and De Iorio (2022).

### v2

@@ 1 Introduction
$f(y\mid\theta_{1},\dots,\theta_{M},\pi)=\sum_{j=1}^{M}\pi_{j}f(y\mid\theta_{j}),$ (1.1)
@@
where $M\in\{1,2,\dots\}$ is the number of components, $y\in\mathbb{R}^{d}$ is a $d$-dimensional observation vector, and $\pi=(\pi_{1},\dots,\pi_{M})^{\top}$ is the mixing weight vector, satisfying $\sum_{j=1}^{M}\pi_{j}=1$ and $\pi_{j}>0$. For $j=1,\dots,M$, $f(\cdot\mid\theta_{j})$ is a probability density or probability mass function parameterized by the component-specific parameter $\theta_{j}$. The function $f(\cdot\mid\theta_{j})$ is also called the kernel of the mixture model. Each observation is assumed to arise from one of the $M$ components, and each component is weighted based on its frequency of occurrence. Mixture models are important statistical models used in model-based clustering and density estimation because they can model complex data-generating distributions in random phenomena by using multiple components. Finite mixture models have been applied in many fields, for instance, sociology (handcock2007model), economics (frohwirth2008model) and genetics (mclachlan2002mixture) and so on. In applications, the determination of $M$ is a very important issue. Correctly estimating $M$ improves model interpretability, the accuracy of kernel parameter estimates and model predictions, and reduces computation time. Although various methods have been proposed for the selection of $M$, such as model selection criteria and hypothesis testing, using these criteria make it difficult to quantify the uncertainty of $M$, and also introduces the bias that model selection is carried out. A comprehensive review of (finite) mixture models is provided by mclachlan2000finite; fruhwirth2006finite; fruhwirth2019handbook; mclachlan2019finite. In this paper, we focus on the case where $M$ is finite and make a clear distinction between the number of components $M$ and the number of clusters $k$. In general, the number of components is written as $M=k+M_{na}$, where $k$ is the number of components for which data are actually assigned, and $M_{na}$ is the number of empty components (see also argiento2022infinity).
Although we also estimate $M$ in this paper, there are two important issues associated with its inference. First, for any finite data set, $M$ is not a well-defined quantity and cannot be directly inferred from the data. In certain application domains of mixture models, such as species sampling in biology (see e.g., argiento2014density), it is even expected that data from some components may not be observed at all. Thus, given the data, we can only make predictions about $M$, and the prior on $M$ and the prior on the weight distribution play a crucial role. Second, the posterior distribution of $M$ is highly sensitive to model misspecification. Most consistency results for $M$ are derived under the assumption of model correctness (nobile1994bayesian; miller2023consistency). Therefore, when attempting to infer the number of mixture components, it is critically important to ensure that the data-generating process is appropriately captured.
@@
In Bayesian analysis, a finite mixture model that assumes prior distributions for the mixing weights and the number of components is usually employed (see, e.g. nobile1994bayesian; miller2018mixture). Such a model is also called a mixture of finite mixtures (MFM), and the model is often used, as well as infinite mixture models represented by Dirichlet process mixture models. Although the reversible jump algorithm (green1995reversible; richardson1997bayesian) has been used to obtain samples from the posterior distribution based on MFM, it faces significant computational and implementation challenges. Methods based on marginal likelihoods have also been proposed (see, e.g. nobile2007bayesian), while, as with reversible jump, the computational aspect is an issue. Recently, the use of nonparametric Bayesian methods for MFM to estimate $M$ has attracted much attention. In miller2018mixture, the exchangeable partition distribution is derived by marginalizing out $M$, and the restaurant process is constructed to overcome computational difficulties. Note that the result of Miller and Harrison (2018) has some connection with Pitman-Yor process prior (see e.g. Section 4.1 in fruhwirth2021generalized. In fruhwirth2021generalized, a generalized MFM in which the parameters of the Dirichlet distribution depend on $M$ is proposed. They showed that this model class may be regarded as a Bayesian nonparametric mixture outside the class of Gibbs-type priors, and proposed a novel sampling algorithm so-called telescoping sampling. In argiento2022infinity, they developed an MFM framework based on a nonparametric Bayesian prior, the normalized independent finite point process. By introducing a gamma latent variable, their approach encompasses a broad class of distributions on the simplex and yields an efficient blocked Gibbs sampling derived from the posterior random measure.
Most studies on MFMs employ a (symmetric) Dirichlet distribution for the mixing weights because of its analytical convenience. argiento2022infinity primarily consider not only the Dirichlet distribution but also a family of simplex distributions constructed from $\sigma$-stable laws. However, the $\sigma$-stable example is treated as a central case, since their paper emphasizes that a restaurant process based computational algorithm can be constructed whenever the Laplace transform of the weight distribution is available.
@@
In this paper, we study a new MFM based on the normalized inverse Gaussian distribution, building on the results of argiento2022infinity. We compare the proposed MFM with the MFM on the Dirichlet prior, focusing on the distributional differences from the Dirichlet case. To the best of our knowledge, there are few studies using the normalized inverse Gaussian distributions as the mixing weights in the field of finite mixture models. In Bayesian nonparametric inference with infinite mixtures, lijoi2005hierarchical proposed a normalized inverse Gaussian process and showed some analytical results of the posterior distribution under the process. Furthermore, we clarify the relationship between the blocked Gibbs sampling and the telescoping sampling, and extend the telescoping sampling so that it can be applied to general distributions on the simplex.
@@
This paper is organized as follows. In Section 2, we provide a brief review of MFMs and its equivalent representation using discrete probability measures. We also discuss the data augmentation and the conditional posterior distribution of $M_{na}$, which are crucial for constructing the computational algorithm. In Section 3, we present the proposal method as well as the posterior computation algorithm. In Section 4, we illustrate some numerical studies to compare the proposed method with existing methods. R code implementing the proposed methods is available at Github repository:
@@ 2 Mixture of finite mixtures
We present a brief overview of a MFM and its equivalent representation of the discrete probability measure. The proposed model and the posterior computation algorithm presented in the next section are largely based on the basic model presented in this section.
@@ 2.1 Formulation
Let each observation $Y_{1},\dots,Y_{n}$ be univariate or multivariate in a Euclidean space. An MFM is a statistical model defined by the following hierarchical representation.
@@
$\begin{split}&M-1\sim q_{M},\quad q_{M}\ \text{is a probability mass function (p.m.f.) on}\ \{0,1,2,\dots\},\\ &\pi\mid M\sim P_{\pi}(\pi\mid M),\\ &c_{i}\mid M,\pi\sim\mathrm{Categorical}_{M}(\pi),\quad i=1,\dots,n,\\ &\tau_{m}\mid M\overset{\mathrm{i.i.d}}{\sim}p_{0}(\tau),\quad m=1,\dots,M,\\ &\theta_{i}\mid c_{i},\tau_{1:M}\overset{\mathrm{ind}}{\sim}\delta_{\tau_{c_{i}}}(\theta_{i}),\quad i=1,\dots n,\\ &Y_{i}\mid\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}\mid\theta_{i}),\quad i=1,\dots,n,\end{split}$ (2.1)
@@
where $f(y\mid\theta_{i})$ is a parametric model with parameter $\theta_{i}$, $\delta_{x}(\cdot)$ is a point mass at $x$ and $\tau_{m}$ is an element of a vector $\tau_{1:M}=(\tau_{1},\dots,\tau_{M})^{\top}$. $p_{0}(\tau_{1:M})$ is a prior density function of the parameter in mixture components. The parameter space is denoted by $\Theta\subset\mathbb{R}^{d}$. Each $c_{i}$ is a latent allocation indicates to which component each observation is allocated. Given the number of components $M$, $P_{\pi}(\pi\mid M)$ is a probability distribution on the $M$-dimensional unit simplex, in that, this is the prior distribution for the mixing weights.
Let $h$ be a probability density function (p.d.f) on $\mathbb{R}_{+}$. We focus on the case where the mixing weights can be further hierarchically expressed as follows:
@@
$\begin{split}&S_{m}\mid M\overset{\mathrm{ind}}{\sim}h,\\ &\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim P_{\pi}(\pi\mid M),\quad T=\sum_{m=1}^{M}S_{m},\end{split}$ (2.2)
where $S_{1},\dots,S_{M}$ are unnormalized weights and $\pi$ is (normalized) mixing weight vector. This representation is the most basic way to generate a probability distribution on the $d$-dimensional unit simplex $\mathbb{S}_{d}:=\{w=(w_{1},\dots,w_{d});w_{i}\geq 0,\sum_{i=1}^{d}w_{i}=1\}$. One of the most famous examples of a distribution on $\mathbb{S}_{d}$ is the Dirichlet distribution. If $S_{i}\sim\text{Gamma}(\gamma_{i},1)$ in (2.2), $\pi\sim\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$, where $\text{Gamma}(a,b)$ is the gamma distribution with shape parameter $a>0$ and rate parameter $b>0$, and $\text{Dirichlet}(\gamma_{1},\dots,\gamma_{d})$ is the Dirichlet distribution with parameter $(\gamma_{1},\dots,\gamma_{d})$. When we set $\gamma_{1}=\cdots=\gamma_{d}\ (=\gamma>0)$, we have the symmetric Dirichlet distribution $\text{Dirichlet}(\gamma,\dots,\gamma)$, and this distribution is often used as a prior distribution for mixing weight vector in the framework of MFM, because of conjugacy with ctegorical distributions and simplicity of computation. The symmetric structure of $\text{Dirichlet}(\gamma,\dots,\gamma)$ is also essential for marginalizing out $M$ and deriving the exchangeable partition probability function (EPPF).
Also, if the hyperparameters of $h$ depend on $M$ in (2.2), the model is referred to as a generalized MFM:
$\begin{split}&S_{m}\mid M\overset{\mathrm{ind}}{\sim}h_{M},\quad m=1,\dots,M,\\ &\pi=\left(\frac{S_{1}}{T},\dots,\frac{S_{M}}{T}\right)\sim P_{\pi}(\pi\mid M),\quad T=\sum_{m=1}^{M}S_{m}.\end{split}$ (2.3)
In particular, if the shape parameters are divided by $M$, the model is referred to as a dynamic MFM (DMFM). The Dirichlet case, where $h_{M}$ is $\text{Gamma}(\gamma/M,1)$, is discussed in fruhwirth2021generalized. On the other hand, fruhwirth2021generalized referred to MFM defined by (2.1), in which the distributional parameters of $S_{m}$ do not depend on $M$, as static MFM. In this paper, we simply refer to such static MFM as MFM.
However, it is known that the estimation result is sensitive to the choice of the shape parameter $\gamma$. For example, miller2018mixture recommended to use $\gamma=1$ as a default choice and the value works well in many cases. On the other hand, in the sparse MFM, dividing by a large $M$ effectively corresponds to using a small value of $\gamma$. Moreover, because $\gamma$ strongly influences both the prior distribution of the number of clusters and the induced partition distribution, it is necessary to assign an appropriate prior to $\gamma$. For further discussion of broader prior specifications, see greve2022spying.
@@ 2.2 Equivalent representations using discrete probability measures
Following argiento2022infinity, we give an equivalent representation of the MFM. We can construct a discrete measure $P(\cdot)=\sum_{m=1}^{M}\pi_{m}\delta_{\tau_{m}}(\cdot)$ in the parameter space $\Theta$, almost surely, where $M$, $\tau$ and $\pi$ are realizations from the distributions $q_{M}$, $p_{0}$ and $P_{\pi}$, respectively. Let $\theta_{1},\dots,\theta_{n}$ be random variables according to $P$. Then, the model (2.1) is equivalent to the following hierarchical representation using a random measure $P$.
$\begin{split}&Y_{i}\mid\theta_{i}\overset{\mathrm{ind}}{\sim}f(y_{i}\mid\theta_{i}),\quad i=1,\dots,n,\\ &\theta_{1},\dots,\theta_{n}\mid P\overset{\mathrm{i.i.d}}{\sim}P,\\ &P\sim\mathcal{P}(q_{M},h,p_{0}),\end{split}$ (2.4)
where $\mathcal{P}$ is a probability distribution of $P$ with parameters $q_{M}$, $h$ and $p_{0}$. The representation (2.4) is the MFM described in the same framework as infinite mixtures. If we replace $\mathcal{P}$ with the Dirichlet process, (2.4) represents the well-known Dirichlet process mixture models (e.g., escobar1995bayesian). If we replace $\mathcal{P}$ with the normalized inverse Gaussian process, (2.4) represents the normalized inverse Gaussian process mixture models (lijoi2005hierarchical). For MFMs, argiento2022infinity proposed a normalized independent finite point process (Norm-IFPP), which is a class of flexible prior distributions for $P$. We employ the representation (2.4) with the Norm-IFPP. The advantages are as follows. We can directly estimate $M$ and $k$. Furthermore, an efficient Gibbs sampler can be constructed by incorporating the data augmentation with a latent Gamma random variable. This data augmentation enables us to overcome the lack of conjugacy in the categorical distribution. Instead of using the probability density function of $P_{\pi}$, we can use the density function of $h$. This is the key to building an efficient MCMC algorithm for the proposed model. The details of Norm-IFPP and the independent finite point process (IFPP) are given in argiento2022infinity.